Silver tarnishes in the presence of hydrogen sulfide (which smells like rotten eggs) and oxygen because of the following reaction: Ag + H2S + O2 --> Ag2S + H2O .

a) How many grams of silver sulfide can be formed from a mixture of 1.1 g Ag, 0.14 g H2S, and 0.080 g O2?



b) How many more grams of hydrogen sulfide would be needed to completely react all the silver?
1
Expert's answer
2013-03-12T12:34:08-0400
4Ag + 2H2S + O2 = 2Ag2S + 2H2O

The amount of substances is

n(mol) = m(g) / MW(g/mol)

n(Ag) = m(Ag) / MW(Ag) = 1.1 / 108 = 0.010mol

n(H2S) = m(H2S) / MW(H2S) = 0.14 / (2 x 1 + 32) =0.0041 mol

n(O2) = m(O2) / MW(O2) = 0.080 / 32 = 0.0025 mol

According to the equation

n(Ag) : n(H2S) : n(O2) = 4: 2 : 1

But actually we have n(Ag): n(H2S) : n(O2) = 4 : 1.64 : 1

As we can see the amount ofH2S is not enough, that’s why H2S is limiting reagent.

n(Ag2S) = n(H2S) = 0.0041mol

The weight of Ag2S is

m(Ag2S) = n(Ag2S) x MW(Ag2S) = 0.0041 x (108 x 2 +32) = 1.017 g

Answer: m(Ag2S) = 1.017 g